# Deep Reinforcement Learning: A Convex Optimization Approach

### v5

@@ 1.1 Background
Deep Reinforcement Learning (RL) has been a cornerstone in most recent developments of Artificial Intelligence One example was defeating the highest ranked player in the ancient game Go (Silver et al., 2017), which in the 90’s was considered a challenge that is hard to crack for the coming 100 years. Another example is the development of chat bots with Large Language Models, based on RL with Human Feedback.
@@
for $\gamma\in(0,1)$, subject to some power constraint on the control signal, $u\in\mathcal{U}$, typically given by a bound on its power, $\|u_{t}\|\leq c_{u}$.
@@ 2 Episodic Deep Reinforcement Learning with Convex Optimization
$Q(x,u,\mathbf{w}^{\star})$ $=c(x,u)+\mathbf{P}_{\mathbf{Q}}\left(\gamma\min_{u_{+}}Q(x_{+},u_{+},\mathbf{w}^{\star})\right)$
@@ 4 Proof of Theorem 1
Using the Bellman equation (LABEL:bellman), we get
@@
For the case where $\alpha_{k}=\alpha$, for all $k\in\mathbb{N}$, we see that the inequality in (12) implies that $w_{k}-w^{\star}$ converges to some $w-w\star$ where

### v6

@@ 1.1 Background
Deep Reinforcement Learning (RL) has been a cornerstone in most recent developments of Artificial Intelligence. One example was defeating the highest ranked player in the ancient game Go (Silver et al., 2017), which in the 90’s was considered a challenge that is hard to crack for the coming 100 years. Another example is the development of chat bots with Large Language Models, based on RL with Human Feedback.
@@
for $\gamma\in(0,1)$, subject to some power constraint on the control signal, $u\in\mathcal{U}$, typically given by a bound on its power, $|u_{t}|\leq c_{u}$.
@@ 2 Episodic Deep Reinforcement Learning with Convex Optimization
$Q(x,u,w^{\star})$ $=c(x,u)+\mathbf{P}_{\mathbf{Q}}\left(\gamma\min_{u_{+}}Q(x_{+},u_{+},w^{\star})\right)$
@@ 4 Proof of Theorem 1
Using the Bellman equation (4), we get
@@
For the case where $\alpha_{k}=\alpha$, for all $k\in\mathbb{N}$, we see that the inequality in (12) implies that $w_{k}-w^{\star}$ converges to some $w-w^{\star}$ where